OpenStudy (anonymous):

Find an equation of the line containing the given point an parallel to the given line. (2,-7);6x-7=4 The equation of the line is Y=

OpenStudy (anonymous):

What is the given line expression?

OpenStudy (anonymous):

6x-7=4 i believe

OpenStudy (anonymous):

6x-7y=4

OpenStudy (anonymous):

Find an equation of the line containing the given point an parallel to the given line. (2,-7);6x-7=4 The equation of the line is Y=

OpenStudy (anonymous):

Solve the given line for y. 6 x - 7 y = 4 -> \[y=\frac{6 }{7}x-\frac{4}{7} \] The new line is:\[y=\frac{6 }{7}x+b \]Plug in the coordinate values of the given point and solve for b:\[-7=\frac{6 }{7}2+b \]\[b=-\frac{61}{7} \]\[y=\frac{6 x}{7}-\frac{61}{7} \] A plot is attached. The red line is the answer and the red dot is the given point.

OpenStudy (anonymous):

thnk u rob. this helps alot